# Conformal Prediction and MLLM aided Uncertainty Quantification in Scene Graph Generation

### v1

@@ Abstract
Scene Graph Generation (SGG) aims to represent visual scenes by identifying objects and their pairwise relationships, providing a structured understanding of image content. However, inherent challenges like long-tailed class distributions and prediction variability necessitate uncertainty quantification in SGG for its practical viability. In this paper, we introduce a novel Conformal Prediction (CP) based framework, adaptive to any existing SGG method, for quantifying their predictive uncertainty by constructing well-calibrated prediction sets over their generated scene graphs. These scene graph prediction sets are designed to achieve statistically rigorous coverage guarantees. Additionally, to ensure these prediction sets contain the most practically interpretable scene graphs, we design an effective MLLM-based post-processing strategy for selecting the most visually and semantically plausible scene graphs within these prediction sets. We show that our proposed approach can produce diverse possible scene graphs from an image, assess the reliability of SGG methods, and improve overall SGG performance.
@@ 1 Introduction
Scene graphs provide a structured semantic representation of a scene depicted in an image. The nodes and edges of this semantic graph structure represent the objects and their bipartite relationships. Such semantic representation of images bridges the gap between vision and language, playing a crucial role in numerous multimodal reasoning tasks such as visual question answering [50], image retreival [18], scene generation [19], and dynamic safety-critical tasks such as autonomous mobility [33] and robot path planning and navigation [37, 1]. However, despite numerous strides made in end-to-end scene graph generation (SGG), this task remains highly challenging due to inherent ambiguities arising from multiple sources, such as imprecise scene descriptions, imperfect object detection, linguistic variations (synonyms and hyponyms) [13, 9, 31, 34], long-tailed class distributions, and multiple viewpoint interpretations. As such, state-of-the-art SGG models tend to produce noisy scene graphs, raising the need to quantify their predictive uncertainty for reliable downstream task applications [24].
@@
While there have been works to incorporate uncertainty minimization in SGG for performance improvement [56], these approaches are designed specifically for their proposed model architectures. In contrast, this paper focuses on a post-hoc distribution-free and model-agnostic uncertainty quantification (UQ) method with formal statistical coverage guarantees. More specifically, we employ Conformal Prediction (CP) [35, 3, 43, 55, 52] to design a model agnostic UQ framework for state-of-the-art SGG models and provide prediction sets of an image’s scene graph. The prediction sets constructed via a CP framework aim to provably contain the true ground-truth class with a desired probability [4]. This provides a method to model the inherent predictive uncertainty of an underlying pre-trained SGG model, with safety assurances [52]. We refer to this form of SGG as Conformal SGG. Fig 1 distinguishes conformal SGG from standard SGG.
@@
Since end-to-end SGG involves detecting all objects in an image and classifying their pairwise relationships, incorporating CP entails the construction of prediction sets for the object and relationship classification. Formally, this pair of objects and their relationship is called a $<subject-predicate-object>$ triplet [23], forming the fundamental block of a scene graph. Therefore, using CP to collectively construct prediction sets for the objects and predicates, we can, in turn, construct a prediction set for each individual triplet. This facilitates the generation of multiple possible scene graphs for an image in the form of prediction sets with formal coverage guarantees
@@
The main objectives of this work are to: 1) design a CP-based post-hoc model-agnostic UQ method adaptive to any existing SGG framework, enabling the generation of multiple possible scene graphs from an image that can provide diverse possible scenarios in a downstream safety-critical task such as robot path planning [37, 36], 2) assess the reliability of existing state-of-the-art SGG methods by analyzing their predictive uncertainty, and 3) show that conformal SGG improves the over recall hit rate compared to standard SGG. As previously mentioned, long-tailed class distribution is a common problem in SGG [13]. Prediction sets from standard CP, which aim to achieve aggregate coverage guarantees across the entire dataset, may fail to provide consistent coverage for each class [40, 14], particularly those in the tail of the distribution. To address this issue, we design an SGG-specific CP framework as class-conditional CP to ensure statistical coverage guarantees at the individual class level as opposed to the dataset level
@@
While CP-based prediction sets of each triplet are constructed to provably achieve desired statistical coverage, depending on the classes captured in the individual object and predicate sets, many entries in the overall triplet prediction set may be visually uncorrelated to the scene and also semantically implausible. This affects the interpretability of the prediction sets and has the practical drawback of large set sizes. To address this issue, we propose a post-processing filtering operation based on scene coherence and feasibility conditions using Multimodal Large Language Models (MLLM) [25]. Having been trained on extensive natural image and language pairs, these models develop an implicit understanding of real-world inter-object relationships, providing a foundational prior for selecting the most plausible scene descriptions associated with an image. Therefore, by converting the entries in each triplet prediction set into language descriptions, we design an effective prompting strategy for in-context learning of an MLLM to achieve the desired task of selecting the most semantically plausible entries in the prediction set without compromising desired coverage guarantees. This operation improves the practicality of the prediction sets by compressing them into the most interpretable detections. We refer to the entire pipeline as Plausibility ensured Conformal SGG, or PC-SGG.
@@
  • We propose a statistically rigorous post-hoc conformal prediction framework for uncertainty quantification in SGG models by generating scene graph prediction sets, which facilitates the generation of diverse scene graphs from an image.
  • To ensure the practical viability of the prediction sets, we propose an MLLM-based post-processing strategy to filter out the set entries that are visually and/or semantically implausible.
  • By using our approach, we assess the reliability of numerous SGG frameworks by analyzing their prediction coverage guarantees.
@@ 2 Related Works
Scene Graph Generation Scene Graph Generation (SGG) predominantly uses detection-based methods, following a two-stage framework: object detection followed by relationship detection. Visual Genome [23] provided foundational insights into the inherent ambiguities in human annotations of visual relationships. Key approaches include message passing structures [29], linguistic information integration [58], region captioning-based methods [30], low-dimensional vector mapping [61], transformer-based networks [48], relational self-attention [26], and feature fusion techniques [12]. Recent incremental methods for relationship detection have emerged in several directions. Some methods use iterative approaches [48, 49] that progressively refine relationship detector parameters, while others focus on predicate representation strategies, including probabilistic modeling [56], fine-grained predicate learning [32], and hybrid-attention mechanisms [15]. Another direction explores sampling strategies, such as bipartite graph networks with adaptive message passing and bi-level data resampling [27]. However, the inherent challenges in SGG [13, 34] necessitate the need for a plug-in tool to assess the predictive uncertainty and, consequently, reliability of these approaches.
@@
Uncertainty in Scene Graph Generation Research on uncertainty in scene graph generation (SGG) has evolved through various approaches. There are works that addressed relationship ambiguity by exploring statistical regularities in scene graphs [60], demonstrating how context influences predicate prediction. Another work [57] introduced an attentional graph convolutional network that implicitly handles uncertainties in relationship proposals. Unbiased scene graph generation is also explored in [47], addressing the uncertainty arising from long-tailed distributions in relationship detection. Moving towards explicit uncertainty modeling,  [17] highlighted semantic ambiguities in visual question answering through scene graphs. However, most of these methods provide strategies for UQ specific to their model architecture and lack safety assurances regarding ground-truth coverage. In this work, we design a model-agnostic UQ method with such statistical guarantees.
@@
Conformal Prediction Conformal Prediction (CP) provides a framework for producing prediction sets with guaranteed coverage under the assumption of exchangeability [54]. The method has gained significant traction in machine learning applications due to its distribution-free nature and rigorous statistical guarantees. In computer vision, CP has been successfully applied to various tasks:  [4] utilized CP for image classification uncertainty quantification, while  [2] extended it to object detection tasks. The framework has also shown promise in causal inference  [51] and time series forecasting [45]. In the context of deep learning, CP has been integrated with neural networks to provide uncertainty estimates in high-stakes applications [22]. However, the application of CP to structured prediction tasks like scene graph generation remains largely unexplored, particularly when dealing with the challenge of large prediction sets while maintaining coverage guarantees.
@@ 3.1 Conformal Prediction
for some tolerated miscoverage rate $\alpha$ [43]. This marginal guarantee provided by generalized CP holds in average across $\mathcal{D}_{test}$ provided $\mathcal{D}_{cal}$ and $\mathcal{D}_{test}$ come from a fixed distribution [52]. In contrast, if granular guarantees are required (for example, at class level or cluster level), conditionally valid coverage is necessary [39, 42, 20].In this work, we are particularly interested in class-conditional coverage [8, 55, 41], yielding the following statistical guarantee, for a test sample $X_{n+1}$,
@@ 4.1 Conformal Scene Graph Generation
Given the ground truth class of the $k^{th}$ triplet is denoted as $y_{k}^{t}=[y_{k}^{s},y_{k}^{r},y_{k}^{o}]\in\mathbb{R}^{3}$ where $y_{k}^{s},y_{k}^{o}\in\mathcal{Y}_{o}$ and $y_{k}^{r}\in\mathcal{Y}_{r}$, the triplet coverage guarantee is given as $P(y_{k}^{t}\in\hat{\mathcal{C}}_{t}(X_{n+1}^{r}))\geq(1-\alpha_{o})(1-\alpha_{r}),\ \forall y_{k}^{s}\in\mathcal{Y}_{o},y_{k}^{o}\in\mathcal{Y}_{o},y_{k}^{r}\in\mathcal{Y}_{r}$.
@@ 4.2 MLLM Guided Plausibility Assessment
While the triplet prediction sets constructed using conformal SGG have nominal coverage guarantees, in practice, many entries in a triplet prediction set may not be plausible. For example, for the image shown under the Inference Scenario in Fig 3, its triplet prediction set has entries like “hand covered in boat, girl holding boat”. The first description of the triplet is implausible from a practical linguistic perspective, while the second one does not correlate to the actual scene in the image and is visually implausible. This causes the overall prediction set of the triplet to be bloated with impractical and uninterpretable scene descriptions. Our proposed PlausiCHECK unit addresses this issue by leveraging the foundation prior of an MLLM in associating the right scene description to the scene at hand. The overview of PlausiCHECK’s functioning is shown in Fig 3.
Plausibility as multiple-choice Q&A. We craft an effective prompting strategy that queries an MLLM regarding the plausibility of an entry in a triplet prediction set. For any detected triplet in a test image, we have the predicted bounding boxes for its subject, $\hat{\mathbf{b}}_{s}$, and object, $\hat{\mathbf{b}}_{o}$, as well as, the prediction set $\hat{\mathcal{C}}_{t}(X)$ obtained via the conformal prediction. First, we convert the entries of $\hat{\mathcal{C}}_{t}(X)$ into language descriptions, then sequentially sample $5$ language descriptions to create a multiple choice question (MCQ), forming the text part of the prompt to the MLLM (Fig 3). The vision part of the prompt is composed of the cropped-out portion of the test image defined by the union box $\hat{\mathbf{b}}_{s}\bigcup\hat{\mathbf{b}}_{o}$. The MLLM is then tasked with answering the MCQ, thereby choosing the most plausible scene description among the $5$ choices. This boils down the next-token prediction of the MLLM into a single-token prediction, with the token space being restricted to the number of choices in the MCQ [44, 16]. In practice, we add a ‘no valid option’ choice to MCQ as shown in Fig 3. In this way, we task the MLLM to eliminate the implausible entries in each $\hat{\mathcal{C}}_{t}(X_{n+1})$ and compress it into a set of practically interpretable scene descriptions.
In order to ensure the MLLM adheres to the task at hand, we leverage a one-shot in-context learning setup [7] and provide adequate task-specific information to the MLLM in the form of an instructional system prompt and a toy scenario of the task at hand in the form of an example prompt (Fig 3). The example prompt contains an image randomly sampled from the calibration set along with a hand-crafted toy MCQA, ensuring the ground-truth triplet is one of the options.
@@ 5.1 Dataset and Splits
Training and Calibration Sets. The calibration set is carved out from the $75$k images in the original VG150150150150 training and validation sets. We design the calibration set in a manner that the images in it cover at least $10\%$ of the total number of samples in each of the object and predicate classes. We ensure that at minimum $2$ samples from any class (object or predicate) are included in the calibration set. The final calibration set contains $7174$ images. The remaining images are used for training.
@@ 5.2 Implementation and Evaluation Metrics
    $Cov_{T}=100\times\frac{1}{n_{t}}\sum\limits_{j=n+1}^{n+n_{t}}m_{j}^{s}\land m_{j}^{r}\land m_{j}^{o}\\ m_{j}^{s}=\mathds{1}[Y_{j}^{s}\in\hat{\mathcal{C}}_{o}(X_{j}^{s})]\\ m_{j}^{r}=\mathds{1}[Y_{j}^{r}\in\hat{\mathcal{C}}_{r}(X_{j}^{r})]\\ m_{j}^{o}=\mathds{1}[Y_{j}^{o}\in\hat{\mathcal{C}}_{o}(X_{j}^{o})]\end{aligned}$
@@
SGG Metrics. The commonly used metrics for quantifying SGG performance are Recall@K (R@K) and mean-Recall@K(mR@K). However, since PC-SGG provides prediction sets for each detected triplet, we modify the computation of the recall hit rate. Instead of checking the equality of the subject, predicate, and object classes in a detected triplet with that of the matched ground truth, we check if the $3$ classes in the matched ground truth are present in the prediction set of the triplet. The remaining aggregation remains the same. We refer to these metrics coverage Recall@K (cR@K) and coverage mean-Recall@K (cmR@K) and can be considered equivalent to R@K and mR@K for evaluating prediction sets of scene graphs.

### v2

@@ Abstract
Scene Graph Generation (SGG) aims to represent visual scenes by identifying objects and their pairwise relationships, providing a structured understanding of image content. However, inherent challenges like long-tailed class distributions and prediction variability necessitate uncertainty quantification in SGG for its practical viability. In this paper, we introduce a novel Conformal Prediction based framework, adaptive to any existing SGG method, for quantifying their predictive uncertainty by constructing well-calibrated prediction sets over their generated scene graphs. These scene graph prediction sets are designed to achieve statistically rigorous coverage guarantees under exchangeability assumptions. Additionally, to ensure the prediction sets contain the most practically interpretable scene graphs, we propose an effective MLLM-based post-processing strategy for selecting the most visually and semantically plausible scene graphs within each set. We show that our proposed approach can produce diverse possible scene graphs from an image, assess the reliability of SGG methods, and improve overall SGG performance.
@@ 1 Introduction
Scene graphs provide a structured semantic representation of a scene depicted in an image. The nodes and edges of this semantic graph structure represent the objects and their bipartite relationships. Such semantic representation of images bridges the gap between vision and language, playing a crucial role in numerous multimodal reasoning tasks such as visual question answering [49], image retreival [18], as well as dynamic safety-critical tasks such as autonomous mobility [32] and robot path planning [36, 1]. However, despite numerous strides made in end-to-end scene graph generation (SGG), this task remains highly challenging due to inherent ambiguities arising from multiple sources, such as imprecise scene descriptions, imperfect object detection, linguistic variations (synonyms and homonyms) [13, 9, 30, 33], long-tailed class distributions, and multiple viewpoint interpretations. As such, state-of-the-art SGG models tend to produce noisy scene graphs, raising the need to quantify their predictive uncertainty for reliable application to downstream tasks [23].
@@
While there have been works to incorporate uncertainty quantification (UQ) in SGG for performance improvement [55], these approaches are designed specifically for their proposed model architectures. In contrast, this paper focuses on a post-hoc distribution-free and model-agnostic UQ method with formal statistical coverage guarantees. Specifically, we employ Conformal Prediction (CP) [34, 3, 54, 51] to design a model agnostic UQ framework for state-of-the-art SGG models and provide prediction sets of an image’s scene graph. The prediction sets constructed via CP aim to provably contain the true ground-truth class with a desired probability [4]. This enables modeling of the inherent predictive uncertainty of an underlying pre-trained SGG model, with safety assurances [51]. We refer to this form of SGG as Conformal SGG, and its distinction with standard SGG is shown in Fig 1.
@@
Since end-to-end SGG involves detecting all objects in an image and classifying their pairwise relationships, incorporating CP entails the construction of prediction sets for the object and relationship classification. Formally, this pair of objects and their relationship is called a $<subject-predicate-object>$ triplet [22], forming the fundamental block of a scene graph. Therefore, by using CP to collectively construct prediction sets for objects and predicates, we can subsequently derive a prediction set for the whole triplet. This facilitates the generation of multiple possible scene graphs for an image in the form of prediction sets with formal coverage guarantees.
@@
The main objectives of this work are to: 1) design a CP-based post-hoc model-agnostic UQ method adaptive to any existing SGG framework, enabling the generation of multiple possible scene graphs from an image that can provide diverse possible scenarios in a downstream safety-critical task such as robot path planning [36, 35], 2) assess the reliability of existing state-of-the-art SGG methods by analyzing their predictive uncertainty, and 3) show that conformal SGG improves the over recall hit rate compared to standard SGG. As previously mentioned, long-tailed class distribution is a common problem in SGG [13], as such prediction sets from standard CP, which aim to achieve aggregate coverage guarantees across the entire dataset, may fail to provide consistent coverage for each class [39, 14], particularly those in the tail of the distribution. To address this issue, we design our SGG-specific CP framework as class-conditional CP [39] to ensure statistical coverage guarantees at the individual class level as opposed to the dataset level.
@@
While CP-based prediction sets of each triplet are constructed to possibly achieve desired statistical coverage, depending on the classes captured in the individual object and predicate sets, many entries in the overall triplet prediction set may be visually uncorrelated to the scene and/or semantically implausible. This affects the interpretability of the prediction sets and has the practical drawback of large set sizes. To address this issue, we propose a post-processing filtering operation based on scene coherence and feasibility conditions using Multimodal Large Language Models (MLLM) [24]. Having been trained on extensive natural image and language pairs, these models develop an implicit understanding of real-world inter-object relationships, providing a foundational prior for selecting the most plausible scene descriptions associated with an image. Therefore, by converting the entries in each triplet prediction set into language descriptions, we design an effective prompting strategy for in-context learning of an MLLM to achieve the desired task of selecting the most semantically plausible entries in the prediction set without compromising desired coverage guarantees. This operation improves the practicality of the prediction sets by compressing them into the most interpretable detections. We refer to the entire pipeline as Plausibility ensured Conformal SGG, or PC-SGG.
@@
  • We propose a statistically rigorous post-hoc conformal prediction framework for UQ of SGG models by generating scene graph prediction sets, which facilitates the generation of diverse scene graphs from an image.
  • To ensure the practical viability of the prediction sets, we propose an MLLM-based post-processing strategy for filtering prediction set entries that are visually and/or semantically implausible.
  • Using PC-SGG, we assess the reliability of numerous SGG frameworks by analyzing the empirical coverage of their prediction sets.
@@ 2 Related Works
Scene Graph Generation. Scene Graph Generation (SGG) predominantly uses detection-based methods, following a two-stage framework: object detection followed by relationship detection. Visual Genome [22] provided foundational insights into the inherent ambiguities in human annotations of visual relationships. Key approaches include message passing structures [28], linguistic information integration [57], region captioning-based methods [29], low-dimensional vector mapping [60], transformer-based networks [47, 25], and feature fusion techniques [12]. Recent methods for relationship detection include iterative approaches [47, 48] that progressively refine relationship detector parameters, while others focus on predicate representation strategies, probabilistic modeling [55], fine-grained predicate learning [31], and hybrid-attention mechanisms [15]. Another direction explores sampling strategies, such as bipartite graph networks with adaptive message passing and bi-level data resampling [26]. However, the inherent challenges in SGG [13, 33] necessitate the need for a plug-in tool to assess the predictive uncertainty and, consequently, reliability of these approaches.
@@
Uncertainty in Scene Graph Generation. Research on uncertainty in scene graph generation (SGG) has largely addressed relationship ambiguity by exploring statistical regularities in scene graphs [59], demonstrating how context influences predicate prediction. Works like [56] introduced an attentional graph convolutional network for implicit uncertainty handling in relationship proposals. Unbiased scene graph generation is also explored in [46], addressing the uncertainty arising from long-tailed distributions in relationship detection. Methods like [17] included uncertainty modeling for highlighting semantic ambiguities in visual question answering through scene graphs. However, most of these methods provide strategies for UQ specific to their model architecture and lack safety assurances regarding ground-truth coverage. In this work, we design a model-agnostic UQ method with such statistical guarantees.
@@
Conformal Prediction. Conformal Prediction (CP) provides a framework for producing prediction sets with guaranteed coverage under the assumption of exchangeability [53]. The method has gained significant traction in machine learning applications due to its distribution-free nature and rigorous statistical guarantees. In computer vision, CP has been successfully applied to various tasks:  [4] utilized CP for image classification uncertainty quantification, while  [2] extended it to object detection tasks. The framework has also shown promise in causal inference  [50] and time series forecasting [44]. In the context of deep learning, CP has been integrated with neural networks to provide uncertainty estimates in high-stakes applications [21]. However, the application of CP to structured prediction tasks like scene graph generation remains largely unexplored, particularly when dealing with the challenge of large prediction sets while maintaining coverage guarantees.
@@ 3.1 Conformal Prediction
for some tolerated miscoverage rate $\alpha$ [42]. This marginal guarantee provided by generalized CP holds in average across $\mathcal{D}_{test}$ provided $\mathcal{D}_{cal}$ and $\mathcal{D}_{test}$ came from a fixed distribution [51]. In contrast, if granular guarantees are required (for example, at class level or cluster level), conditionally valid coverage is necessary [38, 41, 19].In this work, we are particularly interested in class-conditional coverage [8, 54, 40], yielding the following statistical guarantee, for a test sample $X_{n+1}$,
@@ 4.1 Conformal Scene Graph Generation
Given the ground truth class of the $k^{th}$ triplet is denoted as $y_{k}^{t}=[y_{k}^{s},y_{k}^{r},y_{k}^{o}]\in\mathbb{R}^{3}$ where $y_{k}^{s},y_{k}^{o}\in\mathcal{Y}_{o}$ and $y_{k}^{r}\in\mathcal{Y}_{r}$, the triplet coverage guarantee is given as $P(y_{k}^{t}\in\hat{\mathcal{C}}_{t}(X_{n+1}^{r}))=P(Y_{n+1}^{o}\in\hat{\mathcal{C}}_{o}(X_{n+1}^{o})\mid Y_{n+1}^{o}=y_{i}^{o})\cdot P(Y_{n+1}^{r}\in\hat{\mathcal{C}}_{o}(X_{n+1}^{r})\mid Y_{n+1}^{r}=y_{m}^{r})\ \forall\ y_{k}^{s}\in\mathcal{Y}_{o},y_{k}^{o}\in\mathcal{Y}_{o},y_{k}^{r}\in\mathcal{Y}_{r}$.
@@
Corollary 1.
Following Theorem 1, $P(y_{k}^{t}\in\hat{\mathcal{C}}_{t}(X_{n+1}^{r}))\geq(1-\alpha_{o})(1-\alpha_{r}),\ \forall\ y_{k}^{s}\in\mathcal{Y}_{o},y_{k}^{o}\in\mathcal{Y}_{o},y_{k}^{r}\in\mathcal{Y}_{r}$.
@@ 4.2 MLLM Guided Plausibility Assessment
While the triplet prediction sets constructed using conformal SGG have nominal coverage guarantees, in practice, many entries in a triplet prediction set may not be plausible. For example, for the image shown under the Inference Scenario in Fig 3, its triplet prediction set has entries like “hand covered in boat, girl holding boat”. The first description of the triplet is implausible from a practical linguistic/semantic perspective, while the second one does not correlate to the actual scene in the image and is visually implausible. This causes the overall prediction set of the triplet to be bloated with impractical and uninterpretable scene descriptions. We propose a plausibility checking unit which addresses this issue by leveraging the foundational prior of an MLLM in associating the right scene description to the scene at hand. The overview of this unit’s functioning is shown in Fig 3.
Plausibility as multiple-choice Q&A. We craft an effective prompting strategy that query’s an MLLM regarding the plausibility of an entry in a triplet prediction set. For any detected triplet in a test image, we have the predicted bounding boxes for its subject, $\hat{\mathbf{b}}_{s}$, and object, $\hat{\mathbf{b}}_{o}$, as well as, the prediction set $\hat{\mathcal{C}}_{t}(X)$ obtained via the conformal prediction. First, we convert the entries of $\hat{\mathcal{C}}_{t}(X)$ into language descriptions, then sequentially sample $5$ language descriptions to create a multiple choice question (MCQ), forming the text part of the prompt to the MLLM (Fig 3). The vision part of the prompt is composed of the cropped-out portion of the test image defined by the union box $\hat{\mathbf{b}}_{s}\bigcup\hat{\mathbf{b}}_{o}$. The MLLM is then tasked with answering the MCQ, thereby choosing the most plausible scene description among the $5$ choices. This boils down the next-token prediction of the MLLM into a single-token prediction, with the token space being restricted to the number of choices in the MCQ [43, 16]. In practice, we add a ‘no valid option’ choice to MCQ as shown in Fig 3. In this way, we task the MLLM to eliminate the implausible entries in each $\hat{\mathcal{C}}_{t}(X_{n+1})$ and compress it into a set of practically interpretable scene descriptions.
In order to ensure the MLLM adheres to the task at hand, we leverage a one-shot in-context learning setup [7] and provide adequate task-specific information to the MLLM in the form of an instructional system prompt and a toy scenario of the task at hand in the form of an example prompt (Fig 3). The example prompt contains a randomly sampled calibration set image along with a hand-crafted toy MCQA, containing the ground-truth triplet as one of the options.
@@ 5.1 Dataset and Splits
Training and Calibration Sets. The calibration set is carved out from the $75$k images in the original VG150150150150 training and validation sets. The calibration set is designed such that its images cover at most $10\%$ of the total number of samples in each of the object and predicate classes, ensuring that a minimum of $2$ samples from any class is included in the calibration set. The final calibration set contains $7174$ images. The remaining images are used for training.
@@ 5.2 Implementation and Evaluation Metrics
    $Cov_{T}&=100\times\frac{1}{n_{t}}\sum\limits_{j=n+1}^{n+n_{t}}m_{j}^{s}\land m_{j}^{r}\land m_{j}^{o}\\ &m_{j}^{s}=\mathds{1}[Y_{j}^{s}\in\hat{\mathcal{C}}_{o}(X_{j}^{s})],\quad m_{j}^{r}=\mathds{1}[Y_{j}^{r}\in\hat{\mathcal{C}}_{r}(X_{j}^{r})],\\ &m_{j}^{o}=\mathds{1}[Y_{j}^{o}\in\hat{\mathcal{C}}_{o}(X_{j}^{o})]\end{aligned}$
@@
SGG Metrics. The common metrics used for quantifying SGG performance are Recall@K (R@K) and mean-Recall@K(mR@K). However, since PC-SGG provides prediction sets for each detected triplet, we modify the computation of the recall hit rate. Unlike R@K where the equality of the subject, predicate, and object classes in a detected triplet is checked to that of a matched ground truth, we check if the matched ground truth is present in the triplet prediction set obtained from PC-SGG. We refer to these metrics, as coverage-Recall@K (cR@K) and coverage-mean-Recall@K (cmR@K) which are specific to Conformal SGG and can be considered equivalent to standard SGG metrics R@K and mR@K.